OpenStudy (anonymous):

Hi super smart people :D I really need help with this problem, I was sick and missed my Chemistry class yesterday so your help would be highly appreciated! If it takes 50mL of .5 M KOH to completely neutralize 125mL of sulfuric acid solution (H2SO4) what is the concentration of the H2SO4 solution?

OpenStudy (frostbite):

Sense sulfuric acid is a strong inorganic acid we may assume it fully dissociate and we can write the reaction: 2 KOH + H2SO4 -> K2SO4 + 2 H2O From the reaction can we conclude: \[n(H _{2}SO _{4})=\frac{ n(KOH) }{ 2 }\] We can see from the information in the assignment we are able to calculate n(KOH): \[C(KOH)=\frac{ n(KOH) }{ V(KOH) } \rightarrow n(KOH)=C(KOH)*V(KOH)\] In addtion to that we know that: \[C(H _{2}SO _{4})=\frac{ n(H _{2}SO _{4}) }{V(H _{2}SO _{4}) }\] Now you should be able to calculate each part or you can substitute so we get an expression of known variables that would be: \[C(H _{2}SO _{4})=\frac{ \frac{ C(KOH)*V(KOH) }{ 2 } }{ V(H _{2}SO _{4}) }\]

OpenStudy (anonymous):

OMG THANK YOU :D @Frostbite you are awesome ^-^